OpenStudy (anonymous):

the area of the rectangle is 12 square inches. the length is 5 more than wice the width. find the length and width

OpenStudy (anonymous):

So we know that area is Length X Width correct? So lets say the length of the rectangle is 6 and the width is 2. 6 + 5 is 11. Now it's as simple as 11 X 2!

OpenStudy (anonymous):

Hang on I just saw the 5 more then twice the width. Give me one second.

OpenStudy (anonymous):

yeah i have two questions set up like this but dont know where to start

OpenStudy (anonymous):

Okay I got it! 6X2+5=17. 17 is the length and 2 is the width :D

OpenStudy (anonymous):

how did u get that equation 6x2+5=17??

OpenStudy (anonymous):

Because area is length X width correct? you said you have an area of 12. 6 x 2=12. Then you said that the length is 5 more then twice the number of the length. So assuming 6 is your length, 6X2=12+5=17

OpenStudy (anonymous):

awesome thank you so much

OpenStudy (anonymous):

No problem!